OpenStudy (anonymous):

Simplify the expression -4x2(3x − 7).

OpenStudy (anonymous):

28x^2 - 12x^3

OpenStudy (unklerhaukus):

\[-4x^2(3x − 7)=-4x^2\times3x+-4x^2\times-7=-12x^3+28x^2=28x^2-12x^3\]